OpenStudy (anonymous):

Give the slope of each line and then determine whether the two lines are parallel, perpendicular, or neither parallel or perpendicular. 5x+3y=9 20x+12y=0

hero (hero):

Remember, all you need is y = mx + b You have two equations that need to be placed into that form. Once you put both equations in the form y = mx + b, multiply the slopes of both equations together. 1. If the product of the slopes is -1, then you know the lines are perpendicular. 2. If the slopes are the same, then you know the lines are parallel. 3. If neither case exists, then you know that the lines are neither parallel or perpendicular.

OpenStudy (anonymous):

These was the problems I was working out myself. If I remember correctly I have to switch the problems to where they are written in the y=mx+b formula.

hero (hero):

Yup, that's a good start :D

OpenStudy (anonymous):

So for them I will have to divide both sides of the problems in order to get y on the left by itself.

hero (hero):

Yup, maybe we should go to a different platform like vyew.

hero (hero):

That way I can observe what you are doing.

OpenStudy (anonymous):

That's a scary thought lol. I haven't figured out all the little functions yet on this site.

hero (hero):

This is a different site and it will be much easier to use. I'm guessing you've never used a site with online whiteboard.

OpenStudy (anonymous):

My school had an online whiteboard for the tutors, but you had to wait over 4 hours just for them to answer you. Then sometimes they never did.

hero (hero):

Well, that won't happen here.

OpenStudy (anonymous):

Alright.

hero (hero):

Click on that link above and it will take you to the site.